PHYS

A proton is released from rest inside a region of
Answer #1

a1 = E1/q

a2 = E2/q

here net displacement is zero.

s1+s2 = 0

s1 = -s2

0.5*a1*t1^2 = -(v1*t2 - 0.5*a2*t2^2)

0.5*a1*t1^2 = -a1*t1*t2 + 0.5*a2*t2^2

a1*(0.5*t1^2 + t1*t2) = 0.5*a2*t2^2

a2/a1 = (0.5*t1^2 + t1*t2)/(0.5*t2^2)

E2/E1 = (0.5*24.8^2 + 24.8*4.29)/(0.5*4.29^2)

= 44.98
Answer #2

media%2F6a3%2F6a365ac3-ae7a-4e79-871b-34

Answer question